Зх — 2y = 6 4х + Зу = -9
Answer #1

0 3х – 2 = 6 4х+37 -9 2 ( Окт 12 Х– 9y-2 ч 21-6 (=Gх |n/x +9 -27 + -17/ - SI > = S! — 3 - 17 (1) = 3х- 2(-3) = 6 Т+6= 6 - - о